OpenStudy (anonymous):

What is the surface area of two rectangular prisms with a similarity ratio of 2:3?

OpenStudy (anonymous):

Similar solutions to your answer, may help if the ratio of surface areas of prism A and prism B is 3/7, then the ratio of surface areas in terms of the lengths, widths and heights of prims a and b is: [2(la*wa)+2(wa*ha)+2(ha*la)] / [2(lb*wb)+2(wb*hb)+2(hb*lb)] = 3/7 or, [(la*wa)+(wa*ha)+(ha*la)] / [(lb*wb)+(wb*hb)+(hb*lb)] = 3/7 since the prisms are similar, we can substitute la=x*lb, wa=x*lb, and ha=x*hb, where "x" is some scaling factor [(la*wa)+(wa*ha)+(ha*la)] / [(la*wa)*x^2+(wa*ha)*x^2+(ha*la)*x^2] = 3/7 1/x^2 = 3/7, or x = sqrt(7/3) finally, the ratio of the volumes is: [la*wa*ha] / [lb*wb*hb] if we use our substitution, = [la*wa*ha] / [la*x*wa*x*hb*x] = 1/x^3 = (3/7)*sqrt(3/7)